# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v1

@@ 1 Introduction
Most of the existing alignment algorithms are sequential, i.e. they separate the alignment process into a number of successive stages, each utilizing one distinctive category of data. For example, the well-known RLHF approach adopted by (Ouyang et al., 2022) breaks the alignment step into the supervised finetuning (SFT), reward modeling (RM), and reinforcement learning (RL) stages, where the SFT stage uses demonstration data, the RM stage uses the preference data, and the RL stage uses a set of available prompts. Similar strategies have been used in other related works such as (Rafailov et al., 2023; Li et al., 2023; Zhu et al., 2023; Liu et al., 2023). Sequencing of learning tasks may facilitate computation at the expense of inefficient exploitation of data. To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage. This simple example underscores the importance of ensuring the reward model is consistent with the policy model. Such consistency can not be guaranteed with sequential execution of the RM and RL tasks.
The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses. The input to the learning problem, besides these data, also includes a model pre-trained using a large corpus of text data. The output of the learning process is a fine-tuned model aligned with human preferences/values which govern the generation of demonstration and preference data. To integrate these diverse data, we propose a joint learning approach, which is informed by recent advances in Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy and Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021). More specifically, the proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all three kinds of data mentioned above in alignment. Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), and some of their extensions can be seen as particular instances of the proposed formulation. We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF by large margins, especially when the amount of high-quality preference data is relatively limited.
@@ 2.1 Notation
Human Feedback Generation. Assume that the human demonstration $\tau$ is generated from a ground-truth human expert policy $\pi^{\rm E}$, i.e., for each state $s_{t}$, we have $a_{t}\sim\pi^{\rm E}(\cdot|s_{t})$. Moreover, let ‘$\tau_{l}\prec\tau_{w}$’ indicate the winner trajectory $\tau_{w}$ is preferred over the trajectory $\tau_{l}$. Further, we denote $(\tau_{l}\prec\tau_{w})\sim\pi^{\rm P}$ as the preference oracle.
Note that when we present our problem formulation, we will use the above data generation distributions so quantities such as likelihood can be properly defined. Finite-sample version of demonstration and preference data will be introduced subsequently to facilitate algorithm design.
@@ 2.2 Reward Learning from Preference Data
Furthermore, existing work such as Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna and Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences. This greatly reduces the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024). There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022). However, these works often fail when the quality of the preference dataset is low Morimura et al. (2024). In our approach, the inclusion of additional demonstrations can be leveraged to counteract the adverse effects of limited preference quality.
@@ 3.1 Preliminaries
To see the impact on the noisy annotation, consider, for example, the case where a given $s_{t}$, suppose there are two actions $\hat{a}_{t}$ and $\tilde{a}_{t}$, and they are equally good. The SFT data respects the ground truth, and the learned SFT policy says $\pi_{\rm SFT}(\tilde{a}_{t}|s_{t})\approx\pi_{\rm SFT}(\hat{a}_{t}|s_{t})$. However, due to labeling noise, we have a preference data $(s_{t},\hat{a}_{t})\prec(s_{t},\tilde{a}_{t})$, therefore we have learned $R(a_{t},\tilde{a}_{t};\theta)>R(a_{t},\hat{a_{t}};\theta)$, yielding significant difference in the final policy. We refer the readers to Sec. 4 for more discussion.
@@ 3.2 A Meta-Formulation
where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parameterized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parameterized reward model $R(\cdot;\theta)$ to the preference data; $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$; $\alpha\geq 0$ is a balancing coefficient reflecting the relative size of demonstration versus preference data. We refer the above formulation as Alignment with Integrated Human Feedback (AIHF); see Fig. 1 for an illustration of the proposed formulation. The AIHF (8) is a meta-problem that models the alignment problem. It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model. Its components can be customized to yield specific alignment formulations and algorithms. But before diving into various customizations, let us discuss the advantages of this formulation.
@@ 3.3 Specification of AIHF
In this section, we specify the formulation (8). Let us begin with the choice of $L_{1}$. It can be directly instantiated by using an objective similar to (1), which is the likelihood function over the collected expert demonstrations. Note that we aim to optimize the reward parameter $\theta$ to align with human feedback in (8a), thus the objective of $L_{1}$ can be specialized as a maximum likelihood function over expert demonstrations as below:
@@ 3.4 Special Cases of AIHF
We note that now the upper- and lower-level problems are completely decomposable, since the upper-level problem solves for the reward parameterization $\theta$, while the lower-level problem solves for the policy (for the given reward), yielding two separate problems, which are exactly the RM and the RL problems in the typical RLHF approach.
@@ 4 Why AIHF can outperform two-stage alignment approaches?
The first-order condition can be written as:
@@
Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is not negative semi-definite.
@@
That is, the RLHF policy can be seen as the softmax policy for the sum of reward estimators.
@@
$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}.$ (24)
@@
$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}).$
That is, the policy estimate obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences. Slightly abusing notations, let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward. It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}.$
@@
$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}.$
@@
As a second example, again suppose $r(\tau_{1})=r(\tau_{2})$. Consider now the datasets:
{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50.$ (25)
@@
$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}.$
@@ 5 Proposed Algorithm for AIHF Training
Let us remark on the gradient expression of the proposed algorithm. Here, we explain how to construct the gradient estimator $g^{k}_{1}$ for $\nabla_{\theta}L_{1}(\theta_{k})$ defined in (28) and $g^{k}_{2}$ for $\nabla_{\theta}L_{2}(\theta_{k})$ defined in (27). At iteration $k$, to estimate $\nabla_{\theta}L_{1}(\theta_{k})$, we need to sample $N$ expert trajectories $\tau^{\rm E}$ from the expert demonstration dataset $\mathcal{D}^{\rm E}$ and $N$ agent trajectories $\tau^{\prime}$ from the policy $\pi^{k+1}$. Then the gradient estimator $g_{1}^{k}$ can be constructed as: $g_{1}^{k}:=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\theta}R(\tau_{i};\theta)-\nabla_{\theta}R(\tau_{i}^{\prime};\theta)\Big{]}$. Furthermore, through sampling $M$ pairs of preference $(\tau_{l}\prec\tau_{w})$, we can further construct the gradient estimator of $\nabla_{\theta}L_{2}(\theta_{k})$ as: $g_{2}^{k}=\sum_{i=1}^{M}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{i,w};\theta)-R(\tau_{i,l};\theta)\big{)}\Big{)}$. It is interesting to note that, although the demonstration data is used, in the final algorithm, the gradient is always in contrastive form. The difference is that for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
Before we close this section, let us remark on the computational complexity of the proposed algorithm. Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2020).
Specifically, one only needs to perform one (or a few steps) of policy optimization step (such as PPO) to obtain an improved $\pi^{k+1}$ before moving to the next iteration.
As demonstrated in our subsequent numerical results, the computational complexity of our method is comparable or sometimes even lower as compared to their sequential optimization counterparts such as RLHF. In our numerical results, we show that the proposed algorithm AIHF can achieve a comparable performance as RLHF while performing less policy optimization steps.
@@ 6.2 LLM Task
Setting. For the HH dataset, We first fine-tune the language models (Pythia-160M / Pythia-1B) Biderman et al. (2023) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model. Moreover, we supervised fine-tuned the language model using the selected top 10k "chosen" responses and named it the demonstration-SFT model. For the reward model, we use EleutherAI/pythia-1.4B  [^4] https://huggingface.co/EleutherAI/pythia-1.4b for both 160m and 1B policy as its base model.

### v2

@@ 1 Introduction
Most of the existing alignment algorithms are sequential, i.e. they separate the alignment process into a number of successive stages, each utilizing one distinctive category of data. For example, the well-known RLHF approach adopted by (Ouyang et al., 2022) breaks the alignment step into the supervised finetuning (SFT), reward modeling (RM), and reinforcement learning (RL) stages, where the SFT stage uses demonstration data, the RM stage uses the preference data, and the RL stage uses a set of available prompts. Similar strategies have been used in other related works such as (Rafailov et al., 2023; Li et al., 2023; Zhu et al., 2023; Liu et al., 2023). Sequencing of learning tasks may facilitate computation at the expense of inefficient exploitation of data. To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited. In such a situation, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage. This simple example underscores the importance of ensuring the reward model is consistent with the policy model. Such consistency can not be guaranteed with sequential execution of the RM and RL tasks.
The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses. The input to the learning problem, besides these data, also includes a model pre-trained using a large corpus of text data. The output of the learning process is a fine-tuned model aligned with human preferences/values which govern the generation of demonstration and preference data.
To integrate these diverse data, we propose a joint learning approach, which is informed by recent advances in Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy and Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021). More specifically, the proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all three kinds of data mentioned above in alignment. Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), and some of their extensions can be seen as particular instances of the proposed formulation. We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF by large margins, especially when the amount of high-quality preference data is relatively limited.
@@ 2.1 Notation
Human Feedback Generation. Assume that the human demonstration $\tau$ is generated from a ground-truth human expert policy $\pi^{\rm E}$, i.e., for each state $s_{t}$, we have $a_{t}\sim\pi^{\rm E}(\cdot|s_{t})$. Moreover, let ‘$\tau_{l}\prec\tau_{w}$’ indicate the winner trajectory $\tau_{w}$ is preferred over the trajectory $\tau_{l}$. Further, we denote $(\tau_{l}\prec\tau_{w})\sim\pi^{\rm P}$ as the preference oracle. Note that when we present our problem formulation, we will use the above data generation distributions so quantities such as likelihood can be properly defined. Finite-sample version of demonstration and preference data will be introduced subsequently to facilitate algorithm design.
@@ 2.2 Reward Learning from Preference Data
Furthermore, existing work such as Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna and Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences. This greatly reduced the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024). There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022). However, these works often fail when the quality of the preference dataset is low Morimura et al. (2024). In our approach, the inclusion of additional demonstrations can be leveraged to counteract the adverse effects of limited preference quality.
@@ 3.1 Preliminaries
To see the impact on the noisy annotation, consider, for example, the case where a given $s_{t}$, suppose there are two actions $\hat{a}_{t}$ and $\tilde{a}_{t}$, and they are equally good. The SFT data respects the ground truth, and the learned SFT policy says $\pi_{\rm SFT}(\tilde{a}_{t}|s_{t})\approx\pi_{\rm SFT}(\hat{a}_{t}|s_{t})$. However, due to labeling noise, we have a preference data $(s_{t},\hat{a}_{t})\prec(s_{t},\tilde{a}_{t})$, therefore we have learned $R(a_{t},\tilde{a}_{t};\theta)>R(a_{t},\hat{a_{t}};\theta)$, yielding significant difference in the final policy.
@@ 3.2 A Meta-Formulation
where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parametrized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parametrized reward model $R(\cdot;\theta)$ to the preference data and $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$. and $\alpha\geq 0$ is one balancing coefficient reflecting the relative size of demonstration versus preference data. We refer the above formulation as Alignment with Integrated Human Feedback (AIHF); see Fig. 1 for an illustration of the proposed formulation. The AIHF (8) is a meta-problem that models the alignment problem. It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model. Its components can be customized to yield specific alignment formulations and algorithms. But before diving into various customizations, let us discuss the advantages of this formulation.
@@ 3.3 Specification of AIHF
In this section, we specify the formulation (8). Let us begin with the choice of $L_{1}$. It can be directly instantiated by using one objective similar to (1), which is the likelihood function over the collected expert demonstrations. Note that we aim to optimize the reward parameter $\theta$ to align with human feedback in (8a), thus the objective of $L_{1}$ can be specialized as a maximum likelihood function over expert demonstrations as below:
@@ 3.4 Special Cases of AIHF
Noticed that now the upper- and lower-level problems are completely decomposable, since the upper-level problem solves for the reward parameterization $\theta$, while the lower-level problem solves for the policy (for the given reward), yielding two separate problems, which are exactly the RM and the RL problems in the typical RLHF approach.
@@ 4 Why AIHF can outperform two-stage alignment approaches?
The first order condition can be written as:
@@
Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is negative semi-definite.
@@
That is, the RLHF policy can be seen as the softmax policy for the sum of rewards estimator.
@@
$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (24)
@@
$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})$
That is the policy estimated obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences. Abusing notation let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward. It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}$
@@
$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
@@
Consider now the datasets:
{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50$ (25)
@@
$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}$
@@ 5 Proposed Algorithm for AIHF Training
Let us remark on the gradient expression of the proposed algorithm. Here, we explain how to construct the gradient estimator $g^{k}_{1}$ for $\nabla_{\theta}L_{1}(\theta_{k})$ defined in (28) and $g^{k}_{2}$ for $\nabla_{\theta}L_{2}(\theta_{k})$ defined in (27). At iteration $k$, to estimate $\nabla_{\theta}L_{1}(\theta_{k})$, we need to sample $N$ expert trajectories $\tau^{\rm E}$ from the expert demonstration dataset $\mathcal{D}^{\rm E}$ and $N$ agent trajectories $\tau^{\prime}$ from the policy $\pi^{k+1}$. Then the gradient estimator $g_{1}^{k}$ can be constructed as: $g_{1}^{k}:=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\theta}R(\tau_{i};\theta)-\nabla_{\theta}R(\tau_{i}^{\prime};\theta)\Big{]}$. Furthermore, through sampling $M$ pairs of preference $(\tau_{l}\prec\tau_{w})$, we can further construct the gradient estimator of $\nabla_{\theta}L_{2}(\theta_{k})$ as: $g_{2}^{k}=\sum_{i=1}^{M}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{i,w};\theta)-R(\tau_{i,l};\theta)\big{)}\Big{)}$. Here, as a remark, although the demonstration data is used, but in the final algorithm, the gradient is always in constrastive form. The difference is that the for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
Before we close this section, let us remark on the computational complexity of the proposed algorithm. Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2020). Specifically, one only needs to perform one (or a few steps) of policy optimization step (such as PPO) to obtain an improved $\pi^{k+1}$ before moving to the next iteration. As demonstrated in our subsequent numerical results, the computational complexity of our method is comparable or sometimes even lower as compared to their sequential optimization counterparts such as RLHF. In our numerical results, we show that the proposed algorithm AIHF can achieve a comparable performance as RLHF while performing less policy optimization steps.
@@ 6.2 LLM Task
Setting. For the HH dataset, We first fine-tune the language models (Pythia-160M / Pythia-1B) Biderman et al. (2023) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model. Moreover, we also SFT the language model using the selected top 10k chosen responses and name it as demonstration-SFT model. For the reward model, we use EleutherAI/pythia-1.4B  [^4] https://huggingface.co/EleutherAI/pythia-1.4b for both 160m and 1B policy as its base model.